OpenStudy (dmezzullo):

Solve the equation Quantity 7 x minus 17 all over 8 = 4 for x. Please show your work and solution in the space provided.

OpenStudy (dmezzullo):

dont google this either

OpenStudy (dmezzullo):

@ajprincess

OpenStudy (anonymous):

\[\frac{ 7x-17 }{ 8 } = 4\]

OpenStudy (anonymous):

is that the question?

OpenStudy (dmezzullo):

ya

OpenStudy (anonymous):

\[7x-17=4(8)\] \[7x=32+17\]

OpenStudy (anonymous):

as per the prob we have (7x-17)/8 =4 or 7x-17=32 or 7x=32 +17 or x=49/7 0r x=7

OpenStudy (anonymous):

\[7x=49\] \[x=\frac{ 49 }{ 7 }\] \[x = 7\]

OpenStudy (anonymous):

try to see what i did there and tell me if you understand what i did

OpenStudy (dmezzullo):

u didvided 49 by 7 and got 7

OpenStudy (anonymous):

yes. why? isnt that correct?

OpenStudy (dmezzullo):

ya tats correct

OpenStudy (dmezzullo):

ello? anyone wanna help me?

OpenStudy (anonymous):

7x−17=4(8) 7x=32+17 7x=49 x=49/7 x=7

OpenStudy (anonymous):

multiply both sides by 8 7x−17=4(8) add 17 to both sides 7x=32+17 simplify 7x=49 divide both sides by 7 x=49/7 simplify x=7

OpenStudy (dmezzullo):

tyvm